# Fully Dynamic Euclidean $k$-Means

### v2

@@ 17.1 Recourse Before Calling Robustify
Since the epoch lasts for $\ell+1$ updates, the $\left|S_{\textnormal{{init}}}\oplus W^{\prime}\right|$ and $24\ell+$ 2 terms in the right hand side of (66) contributes an amortized recourse of $\tilde{O}(1)$. Moreover, the term $\left|W^{\prime}\oplus S_{\textnormal{{final}}}\right|$ is proportional to the number of MakeRobust calls made while computing $S_{\textnormal{{final}}}\leftarrow\texttt{Robustify}\left(W^{\prime}\right)$. In Lemma 17.2, we show that the amortized number of calls made to the MakeRobust subroutine throughout the entire sequence of updates is at most $\tilde{O}(1)$, which proves Lemma 17.1.
@@
Consider one Type III call $w_{i}\leftarrow\texttt{MakeRobust}(w_{i-1})$. This call is made on 8 in Algorithm 7, which means that $t[w_{i-1}]$ satisfies $t[w_{i-1}]<t$ where $t$ is the smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(w_{i-1},W^{\prime}-w_{i-1})/\lambda^{10}$. Hence,
@@
As a result, $t[w_{0}]<t[w_{1}]<\ldots<t[w_{j}]$. Since the aspect ratio of the space is $(\sqrt{d}\Delta)$, all of the integers $t[w_{i}]$ are between 00 and $\log_{\lambda^{3}}(\sqrt{d}\Delta)<\log(\sqrt{d}\Delta)$ which concludes $j\leq\log(\sqrt{d}\Delta)$. ∎
@@ 18.1 Running Time of Robustify
According to our algorithm, for each center $u$ in the yellow set $\mathcal{Y}$, in $\tilde{O}(1)$ time (5 to 6) [^20] Note that the number $\widehat{\mathrm{dist}}(u,S-u)$ is maintained explicitly by the data structure in Lemma 7.3 ., our algorithm decides whether or not a call to MakeRobust will be made on $u$. Each call to MakeRobust also takes $\tilde{O}(n^{\epsilon})$ time according to Section 18.3. As a result, if the number of elements added to $\mathcal{Y}$ is $N$ and the number of calls to MakeRobust is $M$ throughout the entire algorithm, we conclude that the running time of handling $W_{3}$ is at most $N\cdot\tilde{O}(1)+M\cdot\tilde{O}(n^{\epsilon})$. According to the recourse analysis of our algorithm, the amortized number of calls to MakeRobust is bounded by $M\leq\tilde{O}(1)$.
@@ 18.3 Running Time of MakeRobust
The main for loop in the implementation consists of at most $O(\log_{\lambda}(\sqrt{d}\Delta))$ many iterations. Each iteration is done in $\tilde{O}(n^{\epsilon})$ time according to the data structure (Lemma 7.4) used in 3 of the implementation. It is obvious that all other parts of the algorithm take at most $\tilde{O}(1)$ time. Hence, the total time spent to perform one call to MakeRobust is at most $\tilde{O}(n^{\epsilon})$.
@@ 19 Improving the Update Time to $\tilde{O}(k^{\epsilon})$
In this section, we show how to improve the update time of our algorithm from $\tilde{O}(n^{\epsilon})$ to $\tilde{O}(k^{\epsilon})$. We do this by using dynamic sparsification as a black box, in a similar manner to previous dynamic clustering algorithms [BCG+24, BCF25a]. In [BCF25a], the authors give an algorithm that maintains a $O(1)$-approximation with $\tilde{O}(1)$ recourse and $\tilde{O}(n)$ update time. Using the dynamic $k$-means algorithm of [BCLP23] as a dynamic sparsifier in a black-box manner, they can then improve the update time of their algorithm to $\tilde{O}(k)$, while only incurring $O(1)$ overhead in the approximation and $\tilde{O}(1)$ overhead in the recourse. We speed up the update time of our algorithm from $\tilde{O}(n^{\epsilon})$ to $\tilde{O}(k^{\epsilon})$ using a completely analogous approach, except that we use the sparsifier of [DHS24], which is more well-suited to Euclidean spaces.
@@
The following theorem summarises the relevant properties of the dynamic algorithm of [DHS24], which we call Sparsifier. [^21] The dynamic algorithm of [DHS24] dynamically maintains a coreset of the input space. For the sake of simplicity, we do not define the notion of a coreset and instead focus on the fact that it can be used as a sparsifier. We use this algorithm as a black-box.
Theorem 19.1 ([DHS24]).
@@ 19.2.1 Approximation Ratio Analysis
By Theorem 19.2, at each point in time, we have that $\textsc{cost}(U,S_{i}^{\prime})\leq\alpha\cdot\mathrm{OPT}_{k}(U)$ with probability at least $21-1/\mathrm{poly}(k)\geq 1/$ 2 for each $i\in[L]$. Since each of the algorithms $\mathrm{\textsc{ALG}}_{1}^{\prime},\dots,\mathrm{\textsc{ALG}}_{L}^{\prime}$ uses independent randomness, the probability that $\textsc{cost}(U,S_{i}^{\prime})>\alpha\cdot\mathrm{OPT}_{k}(U)$ for all $i\in[L]$ is at most $(1/2)^{L}=1/\mathrm{poly}(n)$. ∎
@@ 19.2.2 Recourse and Update Time Analysis
Now, let $p(n)=\mathrm{poly}(n)=\Omega(n)$ be the polynomial such that the approximation guarantee of ALG holds with probability at least $1-1/p(n)$ after each update, as described in Theorem 19.2. Consider some subsequence of updates $\sigma_{\ell+1},\dots,\sigma_{\ell+q}$, where $q\leq p(k)/2$. We denote $\sigma_{\ell+j}$ and $r_{\ell+j}$ by $\sigma^{\prime}_{j}$ and $r^{\prime}_{j}$ respectively, and let $r^{\prime}:=\sum_{j=1}^{q}r^{\prime}_{j}$.
@@
We refer to an update $\sigma_{j}^{\prime}$ as bad if $r_{j}^{\prime}=1$, i.e. if we restart the algorithm $\mathrm{\textsc{ALG}}^{\star}$ with fresh randomness at the end of the update. Suppose that $r^{\prime}\geq i-1$ and let $\sigma_{j_{1}}^{\prime},\dots,\sigma_{j_{i-1}}^{\prime}$ be the first $i-1$ bad updates in $\sigma^{\prime}_{1},\dots,\sigma^{\prime}_{q}$. Since we restart $\mathrm{\textsc{ALG}}^{\star}$ after update $\sigma^{\prime}_{j_{i-1}}$, the random bits used by $\mathrm{\textsc{ALG}}^{\star}$ that effect whether any of the updates $\sigma_{j_{i-1}+1}^{\prime},\dots,\sigma_{q}^{\prime}$ are bad are independent of the random bits used by $\mathrm{\textsc{ALG}}^{\star}$ to handle the previous updates. Thus, conditioned on this event, each of the updates $\sigma_{j_{i-1}+1}^{\prime},\dots,\sigma_{q}^{\prime}$ is bad with probability at most $1/p(k)$ by Claim 19.7. Taking a union bound, we get that none of them is bad with probability at least $21-q/p(k)\geq 1/$ 2. ∎
@@
It follows from Lemmas 19.5 and 19.9 that $R=\tilde{O}(T)+k\cdot\tilde{O}(T/k)=\tilde{O}(T)$ with probability at least $1-1/\mathrm{poly}(n)$. Thus, the amortized recourse of our algorithm is $\tilde{O}(1)$ with high probability.
@@
For each $i\in[T]$, let $r^{\star}_{i}$ denote the total number of times that our algorithm resets $\mathrm{\textsc{ALG}}^{\star}$ after handling the update $\sigma_{i}$. Note that $r_{i}\geq r_{i}$ and $r^{\star}=\sum_{i=1}^{T}r_{i}^{\star}$. Now, suppose that we have restarted $\mathrm{\textsc{ALG}}^{\star}$ $j$ times so far after handling the update $\sigma_{i}$, i.e. that $r^{\star}_{i}\geq j$. Since the algorithm uses fresh randomness, the probability that we restart it again (i.e. that $r_{i}^{\star}\geq j+1$) is at most $21/p(k)\leq 1/$ 2. Thus, it follows that
@@
It follows from Lemmas 19.11, 19.10 and 19.9 that the total update time of our algorithm is $\tilde{O}(k^{\epsilon})\cdot T$ with probability at least $1-1/\mathrm{poly}(n)$. Thus, the amortized update time of our algorithm is $\tilde{O}(k^{\epsilon})$ with high probability.
@@ 20 Deferred Proofs
The first inequality holds by Equation 70, the third one holds by Lemma 15.1 (note that $w_{i}^{\prime}=$ MakeRobust$(w_{i})$) and the last one holds by Equation 71 (by exchanging $i$ and $j$ in this equation since $i<j$). Since $\lambda$ is a large number, we conclude that for each $i<j$, the following holds
@@
[…]
@@
Consider the standard LP relaxation for the weighted $k$-means problem. We consider the set of potential centers to open as $S$, and the goal is to open at most $k-(m-r)/4$ many centers, i.e., find $\tilde{S}\subseteq S$.
@@
Each center $u\in S$ that forms a well-separated pair with a center $v\in V$ has $|\pi^{-1}_{S}(u)|\geq 1$ since $u$ must be the closest center to $v$ in $S$. Since the number of well-separated pairs is $k-m$, we have
@@
Finally, since the integrality gap of the LP relaxation is known to be at most $9$ [ANFSW20]. [^23] Note that although it is not explicitly mentioned, the result of [ANFSW20] works for the weighted case. One way to show this, which only uses the unweighted case as a black box is as follows. First, we can round all the real weights to some rational weights with an additive error of at most $\epsilon$. This will not change the feasible solutions and change the optimal objective function only by a factor of $1\pm\epsilon$. Then, we scale up all the weights to have integer weights. Now, we replace every point $p$ of weight $w(p)$ with $w(p)$ many unweighted points $p_{1},p_{2},\cdots,p_{w(p)}$. It is easy to find a correspondence between any optimal solution (either fractional or integral) of the original weighted LP and any optimal solution of the new unweighted LP. Hence, we conclude that the integrality gap of the weighted LP is also at most $9+O(\epsilon)$. Since this is correct for every $\epsilon>0$, we conclude that the integrality gap is at most $9$. As a result, there exist an integral solution whose cost is at most $81\lambda^{40}\left(\textsc{cost}(X,S)+\textsc{cost}(X,V)\right)$, and this solution opens at most $k-\lfloor\frac{m-r}{4}\rfloor$ centers which completes the proof.
@@
Denote the cost of the optimal fractional solution for this LP by $\text{FOPT}_{k}$. Since the space $X$ is fixed here, we denote $\mathrm{OPT}_{k}(X)$ by $\mathrm{OPT}_{k}$. It is known that the integrality gap of this relaxation is at most $9$ [ANFSW20]. So, for every $k$ we have
@@
The first inequality holds since $w^{\prime}$ is present in $W^{\text{new}}$, the second inequality is triangle inequality, the third inequality holds by Equation 84 ,and the last inequality holds since $\lambda=(p_{\epsilon})^{2}$ is a large number. According to 6 in Algorithm 7, we must have $t[w_{0}]<t$, where $t$ is the smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(w_{0},W^{\text{new}}-w_{0})/\lambda^{10}$, as otherwise we should not made a call to MakeRobust on $w_{0}$. Hence,
@@
Appendix A Barriers Towards Adapting the Algorithm of [BCG+24]
@@
In [BCG+24], the authors give a dynamic algorithm for $k$-means on general metric spaces, with $O(1/\epsilon)$-approximation ratio and $\tilde{O}(k^{\epsilon})$ update time. With the current state-of-the-art tools for designing geometric algorithms in Euclidean spaces, we cannot avoid a tradeoff between a $\mathrm{poly}(1/\epsilon)$ term in the approximation and a $k^{\epsilon}$ term in the update time for this problem. [^24] Such tradeoffs are standard in the geometric setting (e.g. from using LSH). Bypassing them, even in the static setting, is a major open problem. Thus, it might appear that building on top of the work of [BCG+24] might be a simpler and more natural approach than building on [BCF25a], since we will eventually incur such a tradeoff in our final algorithm. Unfortunately, the algorithm of [BCG+24] is not as well suited to implementation in high-dimensional Euclidean spaces, and it is not clear how to use this algorithm without avoiding an exponential blowup in the approximation ratio, leading to a $\exp(1/\epsilon)$-approximation.
The Algorithm of [BCG+24].
@@
The high level approach taken by [BCG+24] is to maintain a hierarchy of nested solutions $X\supseteq S_{0}\supseteq\dots\supseteq S_{\ell}$, where $S_{\ell}$ has size $k$ and is the output of the algorithm. Using a new randomized variant of local search, they can maintain this hierarchy by periodically reconstructing these solutions and maintaining them lazily between reconstructions, leading to an update time of $\tilde{O}(k^{1+1/\ell})$ and a recourse of $\tilde{O}(k^{1/\ell})$. To bound their approximation ratio, they show that their randomized local search almost matches the guarantees provided by the projection lemma, which says that there must exist some $S_{i+1}\subseteq S_{i}$ such that $\textsc{cost}(X,S_{i+1})\leq\textsc{cost}(X,S_{i})+O(\mathrm{OPT}_{k}(X))$. Applying this argument $O(\ell)$ times, they show that $\textsc{cost}(X,S_{\ell})\leq O(\ell)\cdot\mathrm{OPT}_{k}(X)$.
The Barrier to Using [BCG+24].
@@
If we want to implement [BCG+24] efficiently in high-dimensional Euclidean spaces, the most natural approach is to modify some components within the algorithm to use approximate nearest neighbor data structures, which can be implemented efficiently in this setting. However, the hierarchical nature of this algorithm requires us to have very strong guarantees on the quality of the solutions in order for the analysis to go through. For example, a guarantee that $\textsc{cost}(X,S_{i+1})\leq O(\textsc{cost}(X,S_{i})+\mathrm{OPT}_{k}(X))$ for the solutions in the hierarchy is not enough to obtain an approximation ratio of $O(\ell)$, and instead would lead to an approximation ratio of $2^{O(\ell)}$ if analyzed in the naive manner. Unfortunately, if we use approximate nearest neighbor data structures to speed up the randomized local search, it is not clear how we can get the sufficiently strong guarantees required to obtain an approximation ratio of $\mathrm{poly}(\ell)$ with this algorithm. To summarize, the lack of robustness to the constants in the guarantees makes the algorithm less well-suited to implementation in high-dimensional Euclidean spaces.

### v3

@@ 18.1 Recourse Before Calling Robustify
Since the epoch lasts for $\ell+1$ updates, the $\left|S_{\textnormal{{init}}}\oplus W^{\prime}\right|$ and $4\ell+2$ terms in the right hand side of (67) contributes an amortized recourse of $\tilde{O}(1)$. Moreover, the term $\left|W^{\prime}\oplus S_{\textnormal{{final}}}\right|$ is proportional to the number of MakeRobust calls made while computing $S_{\textnormal{{final}}}\leftarrow\texttt{Robustify}\left(W^{\prime}\right)$. In Lemma 18.2, we show that the amortized number of calls made to the MakeRobust subroutine throughout the entire sequence of updates is at most $\tilde{O}(1)$, which proves Lemma 18.1.
@@
Consider one Type III call $w_{i}\leftarrow\texttt{MakeRobust}(w_{i-1})$. This call is made on 15 in Algorithm 7, which means that $t[w_{i-1}]$ satisfies $t[w_{i-1}]<t$ where $t$ is the smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(w_{i-1},W^{\prime}-w_{i-1})/\lambda^{10}$. Hence,
@@
As a result, $t[w_{0}]<t[w_{1}]<\ldots<t[w_{j}]$. Since the aspect ratio of the space is $(\sqrt{d}\Delta)$, all of the integers $t[w_{i}]$ are between 0 and $\log_{\lambda^{3}}(\sqrt{d}\Delta)<\log(\sqrt{d}\Delta)$ which concludes $j\leq\log(\sqrt{d}\Delta)$. ∎
@@ 19.1 Running Time of Robustify
According to our algorithm, for each center $u$ in the yellow set $\mathcal{Y}$, in $\tilde{O}(1)$ time (10 to 12) [^20] Note that the number $\widehat{\mathrm{dist}}(u,S-u)$ is maintained explicitly by the data structure in Lemma 8.3 ., our algorithm decides whether or not a call to MakeRobust will be made on $u$. Each call to MakeRobust also takes $\tilde{O}(n^{\epsilon})$ time according to Section 19.3. As a result, if the number of elements added to $\mathcal{Y}$ is $N$ and the number of calls to MakeRobust is $M$ throughout the entire algorithm, we conclude that the running time of handling $W_{3}$ is at most $N\cdot\tilde{O}(1)+M\cdot\tilde{O}(n^{\epsilon})$. According to the recourse analysis of our algorithm, the amortized number of calls to MakeRobust is bounded by $M\leq\tilde{O}(1)$.
@@ 19.3 Running Time of MakeRobust
The main for loop in the implementation consists of at most $O(\log_{\lambda}(\sqrt{d}\Delta))$ many iterations. Each iteration is done in $\tilde{O}(n^{\epsilon})$ time according to the data structure (Lemma 8.4) used in 4 of the implementation. It is obvious that all other parts of the algorithm take at most $\tilde{O}(1)$ time. Hence, the total time spent to perform one call to MakeRobust is at most $\tilde{O}(n^{\epsilon})$.
@@ 20 Improving the Update Time to $\tilde{O}(k^{\epsilon})$
In this section, we show how to improve the update time of our algorithm from $\tilde{O}(n^{\epsilon})$ to $\tilde{O}(k^{\epsilon})$. We do this by using dynamic sparsification as a black box, in a similar manner to previous dynamic clustering algorithms [BhattacharyaCGLP24, BCF24]. In [BCF24], the authors give an algorithm that maintains a $O(1)$-approximation with $\tilde{O}(1)$ recourse and $\tilde{O}(n)$ update time. Using the dynamic $k$-means algorithm of [nips/BhattacharyaCLP23] as a dynamic sparsifier in a black-box manner, they can then improve the update time of their algorithm to $\tilde{O}(k)$, while only incurring $O(1)$ overhead in the approximation and $\tilde{O}(1)$ overhead in the recourse. We speed up the update time of our algorithm from $\tilde{O}(n^{\epsilon})$ to $\tilde{O}(k^{\epsilon})$ using a completely analogous approach, except that we use the sparsifier of [esa/TourHS24], which is more well-suited to Euclidean spaces.
@@
The following theorem summarises the relevant properties of the dynamic algorithm of [esa/TourHS24], which we call Sparsifier. [^21] The dynamic algorithm of [esa/TourHS24] dynamically maintains a coreset of the input space. For the sake of simplicity, we do not define the notion of a coreset and instead focus on the fact that it can be used as a sparsifier. We use this algorithm as a black-box.
Theorem 20.1 ([esa/TourHS24]).
@@ 20.2.1 Approximation Ratio Analysis
By Theorem 20.2, at each point in time, we have that $\textsc{cost}(U,S_{i}^{\prime})\leq\alpha\cdot\mathrm{OPT}_{k}(U)$ with probability at least $1-1/\mathrm{poly}(k)\geq 1/2$ for each $i\in[L]$. Since each of the algorithms $\mathrm{\textsc{ALG}}_{1}^{\prime},\dots,\mathrm{\textsc{ALG}}_{L}^{\prime}$ uses independent randomness, the probability that $\textsc{cost}(U,S_{i}^{\prime})>\alpha\cdot\mathrm{OPT}_{k}(U)$ for all $i\in[L]$ is at most $(1/2)^{L}=1/\mathrm{poly}(n)$. ∎
@@ 20.2.2 Recourse and Update Time Analysis
Now, let $p(n)=\mathrm{poly}(n)=\Omega(n)$ be the polynomial such that the approximation guarantee of ALG holds with probability at least 11/p(n)1-1/p(n) after each update, as described in Theorem 20.2. Consider some subsequence of updates $\sigma_{\ell+1},\dots,\sigma_{\ell+q}$, where $q\leq p(k)/2$. We denote $\sigma_{\ell+j}$ and $r_{\ell+j}$ by $\sigma^{\prime}_{j}$ and $r^{\prime}_{j}$ respectively, and let $r^{\prime}:=\sum_{j=1}^{q}r^{\prime}_{j}$.
@@
We refer to an update $\sigma_{j}^{\prime}$ as bad if $r_{j}^{\prime}=1$, i.e. if we restart the algorithm $\mathrm{\textsc{ALG}}^{\star}$ with fresh randomness at the end of the update. Suppose that $r^{\prime}\geq i-1$ and let $\sigma_{j_{1}}^{\prime},\dots,\sigma_{j_{i-1}}^{\prime}$ be the first i1i-1 bad updates in $\sigma^{\prime}_{1},\dots,\sigma^{\prime}_{q}$. Since we restart $\mathrm{\textsc{ALG}}^{\star}$ after update $\sigma^{\prime}_{j_{i-1}}$, the random bits used by $\mathrm{\textsc{ALG}}^{\star}$ that effect whether any of the updates $\sigma_{j_{i-1}+1}^{\prime},\dots,\sigma_{q}^{\prime}$ are bad are independent of the random bits used by $\mathrm{\textsc{ALG}}^{\star}$ to handle the previous updates. Thus, conditioned on this event, each of the updates $\sigma_{j_{i-1}+1}^{\prime},\dots,\sigma_{q}^{\prime}$ is bad with probability at most $1/p(k)$ by Claim 20.7. Taking a union bound, we get that none of them is bad with probability at least $1-q/p(k)\geq 1/2$. ∎
@@
It follows from Lemmas 20.5 and 20.9 that $R=\tilde{O}(T)+k\cdot\tilde{O}(T/k)=\tilde{O}(T)$ with probability at least $1-1/\mathrm{poly}(n)$. Thus, the amortized recourse of our algorithm is $\tilde{O}(1)$ with high probability.
@@
For each $i\in[T]$, let $r^{\star}_{i}$ denote the total number of times that our algorithm resets $\mathrm{\textsc{ALG}}^{\star}$ after handling the update $\sigma_{i}$. Note that $r_{i}\geq r_{i}$ and $r^{\star}=\sum_{i=1}^{T}r_{i}^{\star}$. Now, suppose that we have restarted $\mathrm{\textsc{ALG}}^{\star}$ $j$ times so far after handling the update $\sigma_{i}$, i.e. that $r^{\star}_{i}\geq j$. Since the algorithm uses fresh randomness, the probability that we restart it again (i.e. that $r_{i}^{\star}\geq j+1$) is at most $1/p(k)\leq 1/2$. Thus, it follows that
@@
It follows from Lemmas 20.11, 20.10 and 20.9 that the total update time of our algorithm is $\tilde{O}(k^{\epsilon})\cdot T$ with probability at least $1-1/\mathrm{poly}(n)$. Thus, the amortized update time of our algorithm is $\tilde{O}(k^{\epsilon})$ with high probability.
@@ 21 Deferred Proofs
[…]
@@
$\textsc{cost}(B_{t^{*}},x_{t^{*}})/w(B_{t^{*}})\leq\lambda^{6t^{*}-2}\text{ and }\textsc{cost}(B_{t^{*}},x_{t^{*}-1})\leq(p_{\epsilon})^{3}\cdot\mathrm{OPT}_{1}(B_{t^{*}}).$
@@
Consider the standard LP relaxation for the weighted $k$-means problem. We consider the set of potential centers to open as $S$, and the goal is to open at most k(mr)/4k-(m-r)/4 many centers, i.e., find $\tilde{S}\subseteq S$.
@@
Each center $u\in S$ that forms a well-separated pair with a center $v\in V$ has $|\pi^{-1}_{S}(u)|\geq 1$ since $u$ must be the closest center to $v$ in $S$. Since the number of well-separated pairs is kmk-m, we have
@@
Finally, since the integrality gap of the LP relaxation is known to be at most $9$ [integrality-gap-k-means]. [^23] Note that although it is not explicitly mentioned, the result of [integrality-gap-k-means] works for the weighted case. One way to show this, which only uses the unweighted case as a black box is as follows. First, we can round all the real weights to some rational weights with an additive error of at most $\epsilon$. This will not change the feasible solutions and change the optimal objective function only by a factor of $1\pm\epsilon$. Then, we scale up all the weights to have integer weights. Now, we replace every point $p$ of weight $w(p)$ with $w(p)$ many unweighted points $p_{1},p_{2},\cdots,p_{w(p)}$. It is easy to find a correspondence between any optimal solution (either fractional or integral) of the original weighted LP and any optimal solution of the new unweighted LP. Hence, we conclude that the integrality gap of the weighted LP is also at most $9+O(\epsilon)$. Since this is correct for every $\epsilon>0$, we conclude that the integrality gap is at most $9$. As a result, there exist an integral solution whose cost is at most $81\lambda^{40}\left(\textsc{cost}(X,S)+\textsc{cost}(X,V)\right)$, and this solution opens at most $k-\lfloor\frac{m-r}{4}\rfloor$ centers which completes the proof.
@@
Denote the cost of the optimal fractional solution for this LP by $\text{FOPT}_{k}$. Since the space $X$ is fixed here, we denote $\mathrm{OPT}_{k}(X)$ by $\mathrm{OPT}_{k}$. It is known that the integrality gap of this relaxation is at most $9$ [integrality-gap-k-means]. So, for every $k$ we have
@@
The first inequality holds since $w^{\prime}$ is present in $W^{\text{new}}$, the second inequality is triangle inequality, the third inequality holds by Equation 85 ,and the last inequality holds since $\lambda=(p_{\epsilon})^{2}$ is a large number. According to 12 in Algorithm 7, we must have $t[w_{0}]<t$, where $t$ is the smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(w_{0},W^{\text{new}}-w_{0})/\lambda^{10}$, as otherwise we should not made a call to MakeRobust on $w_{0}$. Hence,
@@
Appendix A Barriers Towards Adapting the Algorithm of [BhattacharyaCGLP24]
@@
In [BhattacharyaCGLP24], the authors give a dynamic algorithm for $k$-means on general metric spaces, with $O(1/\epsilon)$-approximation ratio and $\tilde{O}(k^{\epsilon})$ update time. With the current state-of-the-art tools for designing geometric algorithms in Euclidean spaces, we cannot avoid a tradeoff between a $\mathrm{poly}(1/\epsilon)$ term in the approximation and a $k^{\epsilon}$ term in the update time for this problem. [^24] Such tradeoffs are standard in the geometric setting (e.g. from using LSH). Bypassing them, even in the static setting, is a major open problem. Thus, it might appear that building on top of the work of [BhattacharyaCGLP24] might be a simpler and more natural approach than building on [BCF24], since we will eventually incur such a tradeoff in our final algorithm. Unfortunately, the algorithm of [BhattacharyaCGLP24] is not as well suited to implementation in high-dimensional Euclidean spaces, and it is not clear how to use this algorithm without avoiding an exponential blowup in the approximation ratio, leading to a $\exp(1/\epsilon)$-approximation.
The Algorithm of [BhattacharyaCGLP24].
@@
The high level approach taken by [BhattacharyaCGLP24] is to maintain a hierarchy of nested solutions $X\supseteq S_{0}\supseteq\dots\supseteq S_{\ell}$, where $S_{\ell}$ has size $k$ and is the output of the algorithm. Using a new randomized variant of local search, they can maintain this hierarchy by periodically reconstructing these solutions and maintaining them lazily between reconstructions, leading to an update time of $\tilde{O}(k^{1+1/\ell})$ and a recourse of $\tilde{O}(k^{1/\ell})$. To bound their approximation ratio, they show that their randomized local search almost matches the guarantees provided by the projection lemma, which says that there must exist some $S_{i+1}\subseteq S_{i}$ such that $\textsc{cost}(X,S_{i+1})\leq\textsc{cost}(X,S_{i})+O(\mathrm{OPT}_{k}(X))$. Applying this argument $O(\ell)$ times, they show that $\textsc{cost}(X,S_{\ell})\leq O(\ell)\cdot\mathrm{OPT}_{k}(X)$.
The Barrier to Using [BhattacharyaCGLP24].
@@
If we want to implement [BhattacharyaCGLP24] efficiently in high-dimensional Euclidean spaces, the most natural approach is to modify some components within the algorithm to use approximate nearest neighbor data structures, which can be implemented efficiently in this setting. However, the hierarchical nature of this algorithm requires us to have very strong guarantees on the quality of the solutions in order for the analysis to go through. For example, a guarantee that $\textsc{cost}(X,S_{i+1})\leq O(\textsc{cost}(X,S_{i})+\mathrm{OPT}_{k}(X))$ for the solutions in the hierarchy is not enough to obtain an approximation ratio of $O(\ell)$, and instead would lead to an approximation ratio of $2^{O(\ell)}$ if analyzed in the naive manner. Unfortunately, if we use approximate nearest neighbor data structures to speed up the randomized local search, it is not clear how we can get the sufficiently strong guarantees required to obtain an approximation ratio of $\mathrm{poly}(\ell)$ with this algorithm. To summarize, the lack of robustness to the constants in the guarantees makes the algorithm less well-suited to implementation in high-dimensional Euclidean spaces.